OpenStudy (anonymous):

2|5x - 2| + 4 = 18 for x.

OpenStudy (anonymous):

i am trying to figure this out...but i can't figure it out.

OpenStudy (anonymous):

is 18x?

OpenStudy (across):

This is basically two equations.

OpenStudy (anonymous):

Solve 2|5x - 2| + 4 = 18 sorry

OpenStudy (across):

\[2(5x-2)+4=18\]and\[2(2-5x)+4=18\]

OpenStudy (anonymous):

Now how did you get that?

OpenStudy (anonymous):

10x-2+4=18 10x-2+4-18 10x-16

OpenStudy (anonymous):

Thank you @Muskan AND @across

OpenStudy (anonymous):

ur wlc

Parth (parthkohli):

A better idea:\[2|5x - 2| + 4 = 18\Rightarrow |5x - 2| = 7\]Square both sides. \[25x^2 - 10x + 4 = 49\]Solve the quadratic equation.